A cube of green jelly rests on a table. The jelly is pushed laterally with a
force of 30 N at it top surface and changes into a parallelogram shape
instead of sliding, with a lateral displacement of 0.125 m. Calculate the
shear strain induced in the cube of jelly
1
Expert's answer
2020-03-16T13:07:03-0400

In the question the angle between the perpendicular y-axis and the top displaced edge of the cube is not mentioned So I am assuming it to be 30o

Here it is given that the lateral displacement of the tip of the vertical edge is 0.125 m. This corresponds to the edge turning through 30 degrees. So we have,

"\tan
30^o
=
\dfrac{0.125}
y
\\y=0.217 m"


Thus the area of a face of the cube is,

"A=y^2=(0.217)^2=0.0468m^2"


The applied force is F=30 N. Therefore the shearing stress is,

"S=\dfrac{F}{A}=\dfrac{30}{0.0468}=641N/m^2"

Shearing strain is "\theta=\pi/6"